# Past, Present, and Future of Citation Practices in HCI

### v1

@@ 5.2. Reflective Considerations for Future Research Practices
Many of the developments in the CHI community are plan-driven. For instance, the number of articles accepted at the CHI conference is deliberately expanded each year. However, one should not forget that top-down policy changes lead to behavioral changes in members of the research community. Notably, while the CHI Steering Committee initially started the CHI community on its expansive path, it is the collective decision of authors at CHI to, year-by-year, cite more and more articles. In this section, we urge the community to pause and reflect on its citation practices.
This work presented trends in citation practices to foster dialogue — not to prescribe mandates, but to encourage diverse interpretations and applications of these findings. We invite readers to engage with the data, reflecting on its implications for the future of HCI research. Our data serves not as a directive, but as a mirror reflecting the current state of CHI’s scholarly practices. By offering this comprehensive overview, we aim to help scholars and practitioners navigate the complexities of the modern research landscape.
However, a critical question arises: How much responsibility does each author in the CHI community bear for the observed changes in citation practices? While scientific knowledge advances rapidly, the academic system remains relatively conservative and resistant to change. This system, dominated by powerful publishers like ACM Press, often feels too vast for any one individual to influence significantly. Therefore, we call for reflective consideration. We urge CHI authors to be more deliberate with their citations, viewing them as quality signals and endorsements. Readers should consider how these trends impact their research practices and the broader HCI community. Through introspection, the community can shape future research in ways that are thoughtful and aligned with shared values. This approach emphasizes the agency of the scholarly community to enact change. It respects the diversity of perspectives within our field and acknowledges each researcher’s autonomy to interpret and integrate these insights according to their goals.
@@ 5.3. Potential Solutions
One potential response to the observed trend is to simply ignore it, though this approach risks future challenges. Proactively addressing the issue of an increasing number of references is a more constructive approach. This section discusses two viable solutions for addressing the growing number of references.
@@ 5.3.1. Solution 1
A first solution would be to revert to the 2015 editorial policy (or an updated version thereof). This could involve setting a limit on the number of references or the overall number of pages per article, as was the case in 2015. At that time, citing a paper was an endorsement, and any citation had to be weighed against the space it would take away from the article’s content or other references. This trade-off has since been lost. While modern practice tends to allow an unrestricted number of pages and references, recent iterations of the CHI Conference have seen inconsistently enforced page limits. For instance, the quality signaling by the Chairs during the CHI ’23 submission process clearly indicated that unrestricted growth in article length is not welcomed by the Conference Chairs and the Steering Committee. Limiting the number of references for most articles (with the exception of literature reviews) would be a viable solution to address the growing and excessive number of references included in CHI articles.
@@ 5.3.2. Solution 2
A second potential solution pertains to the nature of the peer review process in academia. Peer review is the cornerstone of scholarly rigor, serving as a crucial filter that ensures the quality, relevance, and integrity of academic publications. From the perspective of creativity theory (Runco and Jaeger, 2012) and Csikszentmihalyi’s systems model of creativity (Csikszentmihalyi, 2014), the peer review process is a gatekeeping process that validates two critical components of an article: Novelty and usefulness (or appropriateness). As gatekeepers, it is the peer reviewer’s task to assess these two components and, hence, the creativity of a submitted article. In some cases, an extensive related work section may support the evaluation of an article’s novelty, especially in cases where the peer reviewer may not have sufficient knowledge for conducting the review. In an ideal world, a peer reviewer lacking sufficient knowledge of the field would not review an article in the first place. However, in practice, a mismatch between reviewer expertise and with a paper is, unfortunately, rather common (Shah, 2022), especially in large and highly diverse conferences, such as CHI. However, in the role of a gatekeeper, it is the peer reviewer’s duty to be knowledgeable about works in the field. However, in a field as vast and diverse as HCI, this can be challenging.
@@
A potential solution is to split peer review into two parts, one conducted by the human peer reviewer, the other one performed by machine learning. Language models, as knowledge compressors, can assess whether a contribution is original and whether the references are comprehensive more effectively than human reviewers. This would allow human reviewers to focus on assessing the second component — the usefulness and appropriateness of the methodologies, findings, and contributions.
@@
Evidence suggests we are moving towards this direction. In the United States, automated essay scoring systems are already grading student essays (Feathers, 2019; Weatherbed, 2024), and a growing number of peer reviewers use language models to write their reviews. A 2024 study by Liang et al. found that up to 17% of peer reviews at top machine learning conferences are substantially authored by language models (Liang et al., 2024). Liang et al.’s study found an overlap of between 30.9% and 39.2% between the points raised by GPT-4 and human reviewers, and their user study with 308 researchers from 110 institutions in the US highlighted that researchers find AI-generated feedback useful (Liang et al., 2023). Similarly, Sun et al. found automated peer review significantly expedited the review process and improved review coverage, although raising concerns around trust, bias, and agency (Sun et al., 2024).
@@
Automating aspects of peer review could offer significant benefits. The cost of peer review is estimated to be billions of dollars annually (Aczel et al., 2021). In a system strained by an overwhelming volume of literature (Hanson et al., 2023), partial automation can alleviate the burden on reviewers. Further, automating peer review could increase trust in the scientific process by reducing biases, such as the Matthew effect, which has been a concern in academia since 1968 (Merton, 1968). In the future, language models could assist in evaluating novelty, leaving the assessment of usefulness to human reviewers, thereby enhancing the efficiency and reliability of the peer review process.
@@ 6. Conclusion
This article presented meta-research on citation practices at the ACM CHI Conference, the top conference in the field of Human-Computer Interaction. Our segmented regression analysis provided clear evidence of a change of pace in academic working in the CHI community, indicative of broader trends in HCI. We provided empirical evidence of a year-by-year increase in number of references included in CHI articles. The field’s excessive citation practices were enabled by a policy decision that lifted page restrictions on the reference section of CHI articles. The implications of such meso-level policy decisions can be profound. This study highlights the value of meta-research for HCI and the CHI community. More meta-research is needed to carefully consider the broader impact of policy decisions and to provide clear recommendations for stakeholders. It is time for the HCI community to pause and reflect on its citation practices, and constructively discuss solutions. A different pathway forward is needed, in particular with an eye on the research community’s junior members and the sustainability of the HCI field.

### v7

@@ 5.2. Reflective Considerations for Future Research Practices
Many of the developments in the CHI community are plan-driven. For instance, the number of articles accepted at the CHI Conference is deliberately expanded each year. One should not forget that such top-down decisions and policy changes lead to behavioral changes in members of the research community. However, while the CHI Steering Committee initially started the CHI community on its expansive path with its 2016 policy, it is the collective decision of authors at CHI to, year-by-year, cite more and more articles. In this section, we urge the community to pause and reflect on its citation practices.
This work is meant as a zealous vehicle for the members of the CHI research community to open their eyes to the community’s growing problems. We presented trends in citation practices to foster dialogue — not to prescribe mandates, but to encourage diverse interpretations and applications of these findings. We invite readers to engage with the data, reflecting on its implications for the future of HCI research. Our work serves not as a directive, but as a mirror reflecting the current state of CHI’s scholarly practices. We acknowledge that works like ours can be misinterpreted. We would like to explicitly stress that this work is not intended to be a “guide” or “recipe” for writing successful CHI papers. By offering a data-backed overview of growing pain points in the field of HCI, we instead aim to help scholars and practitioners navigate the complexities of the modern research landscape and drive positive change in the community.
However, a critical question arises. How much responsibility does each author in the CHI community bear for the observed changes in citation practices? While scientific knowledge advances rapidly, the academic system remains relatively conservative and resistant to change. This system, dominated by powerful publishers like ACM Press, often feels too vast for any one individual to influence significantly. Therefore, we also call for reflective consideration on an individual level. We urge CHI authors to be more deliberate with their citations, viewing them as quality signals and endorsements. Readers should consider how citation trends impact their research practices and the broader HCI community.
Through introspection, the community can shape future research in ways that are thoughtful and aligned with shared values. The introspective approach emphasizes the agency of the scholarly community to enact change, respecting the diversity of perspectives within our field and acknowledging each researcher’s autonomy to interpret and integrate these insights according to their goals. We should acknowledge that, while fundamental goals (such as publishing articles in high-quality venues) are shared by all of us, the way to achieve these goals varies on an individual level, as demonstrated, for instance, by the ‘slow science’ movement (Alleva, 2006). This movement calls for a more deliberate and cautious approach to publishing. Moving forward, the CHI community should find ways to accommodate these diverse approaches in its citation practices.
@@ 5.3. Potential Solutions
In this section, we discuss potential solutions addressing the observed upward trend in the mean number of references in CHI articles. One potential response is to simply ignore the trend, though this approach risks future challenges. Proactively addressing the issue of excessive citations is a more constructive approach. This section discusses four potential solutions for addressing the growing number of references.
@@ 5.3.1. Solution 1
A first solution would be to revert to the 2015 policy (or an updated version thereof). This could involve setting a limit on the number of references, the number of reference pages, or the overall number of pages per article, as was the case in 2015. At that time, a citation was an endorsement, and any citation had to be weighed against the space it would take away from the article’s content or other references. This trade-off has since been lost.
While modern research practice tends to allow an unrestricted number of pages and references, recent iterations of the CHI Conference have seen strong nudging toward lower page limits. For instance, recent CHI conferences clearly signaled that unrestricted growth in article length is not welcomed by the Conference Chairs and the Steering Committee. The suggested number of words for CHI articles in 2024 was 7000–8000 words – far lower than the average (Oppenlaender and Hämäläinen, 2023). Yet, the strong recommendations are inconsistently enforced by the subcommittees. Limiting the number of references for most articles (with the exception of literature reviews) could be a viable solution to address the growing and excessive number of references included in CHI articles. For such an approach to be successful, hard limits would need to be set, without exceptions, and these limits would need to be consistently enforced by all subcommittees.
@@ 5.3.2. Solution 2
A second solution is to develop language models for supporting paper writing. In particular, specialized language models could be trained to automatically cite prior work, given the text of a manuscript. From the perspective of the author, this would alleviate much of the literature work, and it would help authors concentrate on ideation and conducting high-quality work. However, as the failed release of Meta’s Galactica model demonstrated in 2022 (Heaven, 2022), language models are still prone to hallucinate non-factual and biased information which makes these models, in their present form, not suitable for literature work. This solution would also not alleviate peer reviewer fatigue, which we discuss in the next section.
5.3.3. Solution 3
@@
A third solution pertains to the nature of the peer review process in academia. Peer review is the cornerstone of scholarly rigor, serving as a crucial filter that ensures the quality, relevance, and integrity of academic publications. From the perspective of creativity theory (Runco and Jaeger, 2012) and Csikszentmihalyi’s systems model of creativity (Csikszentmihalyi, 2014), the peer review process is a gatekeeping process that validates two critical components of an article: novelty and importance (also referred to as usefulness or appropriateness) (Runco and Jaeger, 2012). As gatekeeper, it is the peer reviewer’s task to assess these two components and, hence, the creativity of a submitted article. In some cases, an extensive related work section may support the evaluation of an article’s novelty, especially in cases where the peer reviewer may not have sufficient knowledge for conducting the review. But in an ideal world, a peer reviewer lacking sufficient knowledge of the field should not review an article in the first place. In practice, however, a mismatch between reviewer expertise and a paper is, unfortunately, rather common (Shah, 2022), especially in large and highly diverse conferences, such as CHI. In the role of a gatekeeper, it is the peer reviewer’s duty to be knowledgeable about works in the field. In a field as vast and diverse as HCI, this can be challenging.
@@
A potential solution is to split peer review into two parts, one conducted by the human peer reviewer, the other one performed by a machine. Language models are knowledge compressors and can assess whether a contribution is original and whether the references are comprehensive more effectively than human reviewers. This would allow human reviewers to focus on assessing the second component of creativity: the importance of the methodologies, findings, and contributions for the field.
@@
Evidence suggests we are already moving toward this solution. In the United States, automated essay scoring systems are already grading student essays (Weatherbed, 2024) and a growing number of peer reviewers use language models to write their reviews. A study by Liang et al. found that up to 17% of peer reviews at top machine learning conferences are substantially authored by language models (Liang et al., 2024). Liang et al.’s study found an overlap of between 30.9% and 39.2% between the points raised by GPT-4 and human reviewers, and their user study with 308 researchers from 110 institutions in the US highlighted that researchers find AI-generated feedback useful. Similarly, Sun et al. found that automated peer review significantly expedited the review process and improved review coverage, although raising concerns around trust, bias, and agency (Sun et al., 2024).
@@
Automating aspects of peer review could offer significant benefits (Kuznetsov et al., 2024). The cost of peer review is estimated to be billions of dollars annually (Aczel et al., 2021). In a system strained by an overwhelming volume of literature (Hanson et al., 2023), partial automation can alleviate the burden on reviewers. Further, automating peer review could increase trust in the scientific process by reducing biases, such as the Matthew effect, which has been a concern in academia since at least 1968 (Merton, 1968). In the future, language models could assist in evaluating novelty, leaving other aspects to human reviewers, thereby enhancing the efficiency and reliability of the peer review process.
5.3.4. Solution 4
The potential solutions discussed so far have centered around policies and peer review. Another idea focuses on driving change from the author side. In some disciplines, journals require authors to pre-register their studies before conducting them. What if CHI authors pledged to submit their articles with a pre-determined number of references? Authors could set this limit for themselves and they would pre-register this limit publicly before writing and submitting their paper. The pre-registered limit would work as a self-binding contract, forcing authors to keep within their pledged number of references. This – like in the years before the page limit was lifted – would force authors to make a trade-off between references and to think more carefully how many and which references to include in their articles. To incentivize participation in this pre-registration scheme, awards could be given to authors who conform to their pre-registered number of references. The public dataset of pre-registrations could contribute to establishing new community norms and standards for the number of references in CHI articles. However, these standards would not need to be the same for every author. For instance, students could be allowed to include fewer references, acknowledging their role as junior members in the research community. This would alleviate the undue burden of literature work on young scholars.
@@ 6. Conclusion
This article presented meta-research on citation practices at the ACM CHI Conference, the top conference in the field of Human-Computer Interaction. Our exploratory analysis provided clear evidence of a change of pace in academic working in the CHI community, indicative of broader trends in HCI. We provided empirical evidence of a year-by-year increase in the number of references included in CHI articles. The excessive citation practices were enabled by a policy decision that lifted page restrictions on the reference section of CHI articles. This policy decision constitutes a non-linearity in the CHI community’s trajectory and had a destabilizing effect on the community. The implications of this meso-level policy decision are profound, underlining the value of meta-research for the CHI community and HCI. More meta-research is needed to carefully consider the broader impact of policy decisions and to provide clear recommendations for stakeholders. It is time for the HCI community to pause and reflect on its citation practices and constructively discuss solutions. A different pathway forward is urgently needed.
Acknowledgements.
The author aimed to limit this paper to 35 references, but unfortunately, this target was exceeded. I pledge to do better next time.